# Ghost projection via focal-field diffraction catastrophes

### v1

@@ I Introduction
Ghost projection is a method for producing any desired spatial distribution of radiant exposure by superimposing a set of pre-determined illuminating patterns [1, 2, 3, 4]. In essence, the required image is decomposed into a multiple exposure of fixed random patterns. The patterns are assigned different positive weights (e.g. exposure times) which depend on the desired image.
Ghost projection may be viewed as the reverse process to that of classical computational ghost imaging (GI) [5]. GI is a computational imaging technique where a set of known illuminating patterns are incident upon an unknown sample. For each illumination pattern, the corresponding measurements of total intensity transmitted (or reflected) by the sample, referred to as “bucket values,” are recorded by a single-pixel sensor. Neither the set of intensity patterns nor the bucket values alone enable one to determine any positional information about the unknown sample. However, an image of the sample can be reconstructed by exploiting the correlation of the two sets of measurements, i.e., pattern and bucket value pairs [6, 7, 8, 9]. Conversely, in ghost projection, the patterns and desired image are known but the required exposure times/weights (analogous to positive bucket values) are unknown, and must be found.
@@
Figure 1: Schematic of diffraction-catastrophe ghost projection. (a) Parallel lines on the left represent a coherent plane wave traveling to the right, with curved wavefronts indicating the wavefield phase has been modulated (aberrated). The transverse intensity pattern at the focal plane is the “diffraction catastrophe” [10]. (b) Superposing many different focal-plane diffraction catastrophes, for a variety of modulations, is the essence of our technique for “ghost projection” [1, 2, 3, 4]. Note that the “pedestal” is a smooth background that is well-approximated by the ensemble average of the utilized focal-field catastrophes.
To the best of our knowledge, at the time of writing there are four papers on ghost projection [1, 2, 3, 4], all of which employ a mask-based strategy whereby illumination of a variety of spatially random attenuating masks is used to generate a set of spatially random patterns, which can approximate a basis for the space of image projections. [^1] The four indicated papers [1, 2, 3, 4] mainly focus on absorption-only masks, however, mention is also made of how the effects (if present) of weak refraction or diffraction by the masks may be taken into account. These latter effects can be important when the mask feature sizes are sufficiently small, or when the illuminating radiation has a sufficiently high degree of coherence. See, in particular, the calculations based on random-mask Fresnel diffraction in the regime of large Fresnel number, in Ref. [1]. More generally, one can simply measure the ensemble of two-dimensional intensity patterns produced at the plane of imaging—thereby implicitly including whatever refraction or diffraction effects occur—and then employ those patterns as a mathematical basis for the purposes of ghost projection. This latter approach, which is employed in the present paper, has the advantage that it does not need to explicitly invoke a diffraction-physics model of how the intensity and phase at the mask exit surface evolves, upon propagating to the imaging plane downstream of the illuminated mask. If sufficiently many random patterns are chosen, any target image can be well-approximated by a linear combination of patterns with non-negative coefficients.
@@
In this context, consider a coherent collimated beam that is incident upon a focusing lens, thereby producing an Airy disk at the focal plane [11]. If the phase of the beam is modulated non-uniformly before reaching the lens (so that a cross section of the beam has smoothly-varying non-constant phase), then at the focal plane, instead of an Airy disk, we generally observe a highly-structured diffraction pattern (“diffraction catastrophe” [^2] Here and henceforth, the term “catastrophe” is used in the technical sense of “catastrophe theory” [12]. In particular, “diffraction catastrophes” [10] refer to the caustic lines of the geometrical-optics (zero-wavelength) limit, which are softened by the nonzero wavelength so as to be “decorated” with fine diffraction features (Ref. [13], p. 2).) that may be modeled using the “diffraction theory of aberrations” [11]. It is natural to ask whether an ensemble of such diffraction catastrophes—produced by an ensemble of different phase modulations that are sufficiently strong to produce caustics—may be used as a basis for the purposes of ghost projection. This key idea, that we suggested in Ceddia et al. [4], is explored in the present paper.
@@
    For focal-field systems via which a large number of focal-field diffraction catastrophes may be sequentially generated in a reliable and reproducible manner, the method provides an avenue for the dynamic on-demand beam shaping of focused fields. A dynamic sequence of on-demand intensity patterns (or frames), $\{F_{1},F_{2},...\}$, with a given frame rate, can be created by these focal-field diffraction catastrophes. Assuming that approximately $N$ focal-field intensity patterns, $\{I_{1}(\mathbf{r}),I_{2}(\mathbf{r}),\cdots\,I_{N}(\mathbf{r})\}$, can be generated in the time of one frame, then each on-demand intensity pattern in the sequence, $F_{t}$, can be generated via the linear combination
    $F_{t}(\mathbf{r}):=\sum_{k=1}^{N}\tau_{k}I_{k}(\mathbf{r}),$ (1)
    where $\tau_{k}$ represents the associated exposure time [^3] Stated more precisely, $\tau_{k}$ is a dimensionless non-negative weighting coefficient that is proportional to an exposure time. of each focal-field, $I_{k}$. Here and henceforth, $\mathbf{r}$ denotes transverse coordinates in the focal plane (see Fig. 1).
@@
    Ghost projection using diffraction catastrophes circumvents the issue of imperfect focusing elements: so long as we include these imperfections when calculating our diffraction catastrophes, the weight-generating algorithm will naturally generate a different set of exposure times to compensate, and the image reconstruction quality will be unaffected. This requires either that the aberrations (imperfections) are sufficiently well-characterized such that the simulated diffraction catastrophes are accurate, or that the diffraction catastrophes are measured empirically. For instance, if one desires to project a single spot, this form of ghost projection can be used as an indirect form of aberration correction, producing a non-distorted spot at the cost of a pedestal.
@@
    Given the capacity for miniaturization inherent in a focused-beam geometry, the method might also be of utility for lithography, e.g. using aberrated electron-beam foci [14] or aberrated x-ray-beam foci [15].
@@ II Principles of ghost projection using a speckle-field basis
Let $\{I_{k}(\mathbf{r})\}_{k\in\{1,\cdots,M\}}$ define a set of $M$ spatially-random images, such as a statistical ensemble
$\{I_{1}(\mathbf{r}),I_{2}(\mathbf{r}),\cdots,I_{M}(\mathbf{r})\}:=\{I_{k}(\mathbf{r})\},\quad k=1,2,\cdots,M$ (2)
of speckle maps [16] that each have the same correlation length [^4] Here, the “correlation length,” $\ell$, of the speckle map refers to the width of the autocovariance function, for the intensity distribution associated with any particular speckle map. The assumption of statistical spatial stationarity implies $\ell$ is independent of position, for any one speckle map. We also implicitly assume ergodicity, so that spatial averages over one speckle map may be replaced with ensemble averages over the set of all such speckle maps, considered as a stochastic process. $\ell$ which is independent of position. As suggested by the sketch in Fig. 2, this set has many of the characteristics that one would expect of a complete basis. In particular, if we let $\overline{I_{k}}(\mathbf{r})$ denote the ensemble-averaged intensity, which is by assumption independent of position, then the set $\{I_{k}(\mathbf{r})-\overline{I_{k}}(\mathbf{r})\}$ will be near-orthogonal (in the sense that the inner product of any two different patterns is expected to be small relative to the inner product of a pattern with itself). Accordingly, such a set can be employed for the synthesis of functions.
@@
Here the angular brackets denote ensemble average and $\delta(\mathbf{r})$ is the Dirac delta. In our case, rather than the preceding “ideal” completeness relation, we instead have the smoothed completeness relation [18, 19, 1]
$\langle[I_{k}(\mathbf{r})-\overline{I_{k}}][I_{k}(\mathbf{r^{\prime}})-\overline{I_{k}}]\rangle=\mathcal{N}~{}PSF(\mathbf{r}-\mathbf{r}^{\prime}).$ (4)
@@
Here, $PSF(\mathbf{r})$ is a rotationally symmetric point-spread function with characteristic width $\ell$, and $\mathcal{N}$ is a normalization constant that enforces the demand that the area under the point-spread function be equal to unity. From a stochastic perspective, Eq. (4) merely makes the trivial statement that the autocovariance of the background-subtracted speckle field decays to zero over a spatial scale on the order of the speckle size.
@@
Figure 2: Illustration of the ghost-projection concept, for the case of spatially-random speckle maps, in the form given by Eq. (7). Here, the smoothed target pattern (the smiley) is $S(\mathbf{r})\otimes PSF(\mathbf{r})$, the pedestal (or offset) is $\overline{I_{k}}\,\,\overline{B_{k}}$, and each speckle pattern $I_{k}(\mathbf{r})$ is weighted by the positive coefficient $B_{k}$.
@@
From an image-synthesis perspective that encapsulates the ghost-projection concept, let $S(\mathbf{r})$ be an image that one wishes to synthesize. Upon multiplying both sides of Eq. (4) by $S(\mathbf{r})$, integrating over $\mathbf{r}$ and using $\otimes$ to denote two-dimensional convolution, we obtain
$S(\mathbf{r})\otimes PSF(\mathbf{r})=\langle(B_{k}-\overline{B_{k}})[I_{k}(\mathbf{r})-\overline{I_{k}}]\rangle.$ (5)
@@
Here, the decomposition coefficients $B_{k}$—or “bucket values”, in ghost-imaging language [7, 6]—are defined via the inner product
[…]
@@
The right side of Eq. (5) shows how one may form a suitable linear combination of the previously mentioned speckle maps, to form a smoothed version of $S(\mathbf{r})$ that is convolved with the point-spread function (smeared Dirac delta) associated with the smoothed completeness relation. This image synthesis may also be written as [1]
$[S(\mathbf{r})\otimes PSF(\mathbf{r})]+\overline{I_{k}}\,\,\overline{B_{k}}=\langle B_{k}I_{k}(\mathbf{r})\rangle.$ (7)
@@
This result is illustrated in Fig. 2. It may be spoken of as “building signals out of noise”, in the sense that it expresses the smoothed target pattern, $S(\mathbf{r})\otimes PSF(\mathbf{r})$, plus the offset (pedestal), $\overline{I_{k}}\,\,\overline{B_{k}}$, via a linear combination of speckle (noise) maps, $\langle B_{k}I_{k}(\mathbf{r})\rangle$. While the linear combination will in general involve averaging over the infinitely many members in the speckle-field basis, this quantity may be approximated using a finite subset drawn from this overcomplete [20] set.
@@ III Diffraction catastrophes as a random basis for ghost projection
If we now replace the ensemble of speckle maps with an ensemble of focal-field diffraction catastrophes [10] associated with coherent randomly-aberrated collapsing spherical waves, [^5] If the aberrations associated with the distorted focal field are sufficiently strong, the resulting focal-plane caustics constitute the “diffraction catastrophes” [10, 13] referred to in this sentence from the main text. If the aberrations are sufficiently weak—while also being large enough to split a focal spot into a relatively large number of local intensity maxima—the peaked focal-field caustic intensity structure will be be significantly softened. In this latter case, one will instead have distorted arrays of diffraction spots such as those illustrated on p. 541 of Born and Wolf [11]. The methods of our paper may also be applied to such a case, namely where the aberrated focal-field patterns are highly structured, but the aberrations are not sufficiently strong for focal-plane caustics to be evident. A similar statement holds for a complementary limiting case, where the geometrical-optics limit implies that caustic-laden focal-field intensities may be described using ray optics. the scenario in Fig. 2 is modified to that shown in Fig. 1b. The key changes are that the correlation length $\ell$ is now a function of transverse position $\mathbf{r}$ and the pedestal, $\mathcal{P}(\mathbf{r})$, is no longer spatially uniform. Equation (7) thereby generalizes to
[…]
@@
whereby the space-invariant point-spread function has been replaced with a space-variant smearing function (scaled autocovariance) that converts the convolution into a linear integral transform. In practice, this position dependence of the smearing function will often be fairly weak, if one restricts the field of view for the desired ghost projection to be in a sufficiently small vicinity of the optical axis.
@@
From this point we will utilize a vector-space description of pixellated images to lay the foundation for understanding how ghost projection functions in practice, and how it can be optimized. Notions of superposition and weighting of patterns naturally lead to treating the illuminating patterns as elements of a vector space [21], a formalism which is amenable to simulation and creating weight-generating algorithms. In ghost projection, the set of illuminating patterns can be considered a fixed set of function-space vectors, and the weight-generating aspect of ghost projection is analogous to finding a linear combination of these vectors that approximates some desired vector (image), with the constraint that the coefficients be non-negative because negative exposure times are non-physical. A consequence of this non-negativity constraint is that the image can only be synthesized up to the previously mentioned offset that we refer to as the “pedestal” [1, 2].
@@
Let the desired ghost-projection image be $S$, and the set of weights for the illuminating patterns be $\{w_{k}\}_{k\in\{1,\cdots,n\}}$, where each $w_{k}\geq 0$ and each $I_{k}$ has only non-negative entries. Note that for ease of notation, we have suppressed the functional dependence on $\mathbf{r}$. The goal of the computational side of ghost projection is to find some optimum set of weights, $w_{k}$, such that in combination they provide an accurate synthesis of $S$ [1, 2, 3], i.e.,
$\sum_{k=1}^{n}w_{k}I_{k}\approx S+\mathcal{P}.$ (9)
@@
    A significantly more efficient method of synthesizing the desired image, $S$, is to use a suitable iterative numerical method such as the non-negative least squares (NNLS) algorithm [2, 3]. Note that the weighting coefficients $w_{k}$ cannot be negative since they are proportional to the exposure times for each element of the random basis. NNLS can be employed to determine the set $\{w_{k}\geq 0\}$ that minimizes the cost function
    $\mathcal{C}=\left\lVert\sum_{k=0}^{n}w_{k}I_{k}-S\right\rVert_{2},$ (12)
@@
While approach (ii) enforces sparsity in an explicit manner, approach (iii) implicitly imposes sparsity since many of the $w_{k}$ coefficients that are thereby obtained will be very small in magnitude in comparison to the non-negligible coefficients. Given that only a very small fraction of the overcomplete random basis [23] need actually be employed, we only retain those weighting coefficients $w_{k}$ that exceed a specified threshold. The larger the number of members in the random basis from which one can choose, the more efficient the ghost-projection process can become. [^6] We have formalized several schemes to select the sparse set of non-zero pattern weightings to produce a desired image, given an initial large set of basis patterns. These schemes will be described, later in the paper.

### v2

@@ I Introduction
Ghost projection is a method for producing any desired spatial distribution of radiant exposure by superimposing a set of pre-determined illuminating patterns [1, 2, 3, 4, 5]. In essence, the required image is decomposed into a multiple exposure of fixed random patterns. The patterns are assigned different positive weights (e.g. exposure times) which depend on the desired image. Ghost projection may be viewed as the reverse process to that of classical computational ghost imaging [6], from which it inherits its name.
To the best of our knowledge, at the time of writing there are five papers on ghost projection [1, 2, 3, 4, 5]. All of these publications employ a mask-based strategy whereby a variety of spatially-random attenuating masks are illuminated to generate a set of spatially-random patterns, which can approximate a basis for the space of image projections. However, at small detail scales, attenuating masks suffer from diffraction effects that distort the patterns, impeding their effectiveness in ghost projection.
@@
Figure 1: Schematic of diffraction-catastrophe ghost projection. (a) Parallel lines on the left represent a coherent plane wave traveling to the right, with curved wavefronts indicating the wavefield phase has been modulated (aberrated). The transverse intensity pattern at the focal plane is the “diffraction catastrophe” [7]. (b) Superposing many different focal-plane diffraction catastrophes, for a variety of modulations, is the essence of our technique for “ghost projection” synthesis of an arbitrary target image [1, 2, 3, 4, 5]. Note that the “pedestal” is a smooth background that is well approximated by the ensemble average of the utilized focal-field catastrophes.
@@
In this paper, we propose a pattern generation procedure that forgoes attenuating masks and exploits diffraction, rather than being limited by it. In this context, consider a coherent collimated beam that is incident upon a focusing lens, thereby producing an Airy disk at the focal plane [8]. If the phase of the beam is modulated non-uniformly before reaching the lens (so that a cross section of the beam has smoothly-varying non-constant phase), then we generally observe a highly-structured diffraction pattern (“diffraction catastrophe” [^1] Here and henceforth, the term “catastrophe” is used in the technical sense of “catastrophe theory” [9]. In particular, “diffraction catastrophes” [7] refer to the caustic lines of the geometrical-optics (zero-wavelength) limit, which are softened by the nonzero wavelength so as to be “decorated” with fine diffraction features (Ref. [10], p. 2).) at the focal plane instead of an Airy disk (see Fig. 1(a)). The envisioned advantage of this method is that the phase function may be undetailed and slowly varying (aiding ease of fabrication) and yet produce a highly detailed diffraction pattern. It is natural to ask whether an ensemble of such diffraction catastrophes—produced by an ensemble of different phase modulations that are sufficiently strong to produce caustics—may be used as a basis for the purposes of ghost projection. This key idea, suggested in Ceddia et al. [4], is explored in the present paper.
@@
    For systems in which many focal-field diffraction catastrophes may be sequentially generated in a reproducible manner, the method provides an avenue for the dynamic on-demand beam shaping of focused fields. A dynamic sequence of on-demand intensity patterns (or frames), $\{F_{1},F_{2},$...\}, with a given frame rate, can be created by these focal-field diffraction catastrophes. Assuming that approximately $\mathcal{N}$ focal-field intensity patterns, $\{I_{1}(\mathbf{r}),I_{2}(\mathbf{r}),\cdots\,I_{\mathcal{N}}(\mathbf{r})\}$, can be generated in the time of one frame, then each on-demand intensity pattern in the sequence, $F_{t}$, can be generated via the linear combination
    $F_{t}(\mathbf{r}):=\sum_{k=1}^{\mathcal{N}}\tau_{k}I_{k}(\mathbf{r}),$ (1)
    where $\tau_{k}$ represents the associated exposure time [^2] Stated more precisely, $\tau_{k}$ is a dimensionless non-negative weighting coefficient that is proportional to an exposure time. of each illuminating pattern, $I_{k}$. Here and henceforth, $\mathbf{r}$ denotes transverse coordinates in the focal plane (see Fig. 1(a)).
@@
    Ghost projection using diffraction catastrophes circumvents the issue of imperfect focusing elements: so long as we include these imperfections when calculating our diffraction catastrophes, the weight-generating algorithm will naturally generate a different set of exposure times to compensate, and the image reconstruction quality will be unaffected. This requires either that the aberrations (imperfections) are sufficiently well-characterized such that the simulated diffraction catastrophes are accurate, or that the diffraction catastrophes are measured empirically. For instance, if one desires to project a single spot, this form of ghost projection can be used as an indirect form of aberration correction, producing a non-distorted spot at the cost of the “pedestal” mentioned in the caption of Fig. 1.
@@
    Given the capacity for miniaturization inherent in a focused-beam geometry, the method might also be useful for lithography, e.g. using aberrated electron-beam foci [11] or aberrated x-ray-beam foci [12, 13].
@@
As additional context, it is clarifying to recall that classical ghost imaging is a computational imaging technique where a set of known illuminating patterns are incident upon an unknown sample. For each illumination pattern, the corresponding measurements of total intensity transmitted (or reflected) by the sample, referred to as “bucket values,” are recorded by a single-pixel sensor. Neither the set of intensity patterns nor the bucket values alone enable one to determine any positional information about the unknown sample. However, an image of the sample can be reconstructed by exploiting the correlation of the two sets of measurements, i.e., pattern and bucket value pairs [14, 15, 16, 17]. Conversely, in ghost projection [1, 2, 3, 4, 5], the patterns and desired image are known but the required exposure times/weights (analogous to positive bucket values) are unknown, and must be found.
@@ II Principles of ghost projection using a speckle-field basis
In this section we outline the principles of ghost projection as established in Paganin [1] and Ceddia et al. [2, 3]. Let $\{I_{k}(\mathbf{r})\}_{k\in\{1,\cdots,\mathcal{N}\}}$ define a set of $\mathcal{N}$ spatially-random intensity images, such as a statistical ensemble
$\{I_{1}(\mathbf{r}),I_{2}(\mathbf{r}),\cdots,I_{\mathcal{N}}(\mathbf{r})\}:=\{I_{k}(\mathbf{r})\},\quad k=1,2,\cdots,\mathcal{N}$ (2)
of speckle maps [18] that each have the same correlation length [^3] Here, the “correlation length,” $\ell$, of the speckle map refers to the width of the autocovariance function, for the intensity distribution associated with any particular speckle map. The assumption of statistical spatial stationarity implies $\ell$ is independent of position, for any one speckle map. We also implicitly assume ergodicity, so that spatial averages over one speckle map may be replaced with ensemble averages over the set of all such speckle maps, considered as a stochastic process. $\ell$ which is independent of position. As suggested by the sketch in Fig. 2, for sufficiently large $\mathcal{N}$ this set has many of the characteristics that one would expect of a complete basis. In particular, if we let $\overline{I_{k}}\in\mathbb{R}^{+}$ denote the spatially-averaged intensity of the $k\text{th}$ intensity pattern, then the set $\{I_{k}(\mathbf{r})-\overline{I_{k}}\}$ will be near-orthogonal (in the sense that the inner product of any two different patterns is expected to be small relative to the inner product of a pattern with itself). Accordingly, such a set can be employed for the synthesis of functions.
@@
Here the angular brackets denote ensemble average over $k\in\{1,\cdots,\mathcal{N}\}$ and $\delta(\mathbf{r})$ is the Dirac delta. In our case, rather than the preceding “ideal” completeness relation, we instead have the smoothed completeness relation [20, 21, 1]
$\langle[I_{k}(\mathbf{r})-\overline{I_{k}}][I_{k}(\mathbf{r^{\prime}})-\overline{I_{k}}]\rangle=\Upsilon~PSF(\mathbf{r}-\mathbf{r}^{\prime}).$ (4)
@@
Here, $PSF(\mathbf{r})$ is a rotationally symmetric point-spread function with characteristic width $\ell$, and $\Upsilon$ is a normalization constant that enforces the demand that the area under the point-spread function be equal to unity. From a stochastic perspective, Eq. (4) states that the autocovariance of the background-subtracted speckle-field intensity decays to zero over a spatial scale on the order of the speckle size.
@@
Figure 2: Illustration of the ghost-projection concept, for the case of spatially-random speckle maps, in the form given by Eq. (8). Here, the smoothed target pattern (the smiling face) is $S(\mathbf{r})\otimes PSF(\mathbf{r})$, the pedestal (or offset) is $\langle B_{k}\rangle\overline{I}$, and each speckle pattern $I_{k}(\mathbf{r})$ is weighted by the positive coefficient $B_{k}$.
@@
From an image-synthesis perspective that encapsulates the ghost-projection concept, let $S(\mathbf{r})$ be an image that one wishes to synthesize. Upon multiplying both sides of Eq. (4) by $S(\mathbf{r}^{\prime})$, integrating over $\mathbf{r}^{\prime}$ and using $\otimes$ to denote two-dimensional convolution, we obtain
$S(\mathbf{r})\otimes PSF(\mathbf{r})=\langle(B_{k}-\overline{I_{k}}\>\overline{S\>})[I_{k}(\mathbf{r})-\overline{I_{k}}]\rangle.$ (5)
@@
Here, the decomposition coefficients $B_{k}$—or “bucket values”, in ghost-imaging language [15, 14]—are defined via the inner product (dot product)
$B_{k}:=\frac{1}{\Upsilon}\iint I_{k}(\mathbf{r})S(\mathbf{r})\,d\mathbf{r}=:I_{k}\cdot S$ (6)
and
$\overline{S}:=\frac{1}{\Upsilon}\iint S(\mathbf{r})\,d\mathbf{r}.$ (7)
@@
The right side of Eq. (5) shows how one may form a suitable linear combination of the previously mentioned speckle maps, to give a smoothed version of $S(\mathbf{r})$ that is convolved with the point-spread function (smeared Dirac delta) associated with the smoothed completeness relation. The speckle patterns $I_{k}(\mathbf{r})$ have the following properties: the spatial average $\overline{I_{k}}$ is approximately constant with respect to $k$ (to indicate that it is independent of $k$, we shall write this number as $\bar{I}$), and the ensemble average $\langle I(\mathbf{r})\rangle$ is approximately uniform with respect to $\mathbf{r}$, equal to $\bar{I}$ everywhere. Using these assumptions and expanding Eq. (5), we obtain
$S(\mathbf{r})\otimes PSF(\mathbf{r})]+\langle B_{k}\rangle\bar{I}=\left\langle B_{k}I_{k}(\mathbf{r})\right\rangle.$ (8)
@@
This result is illustrated in Fig. 2. It may be spoken of as “building signals out of noise” [1], in the sense that it expresses the smoothed target pattern, $S(\mathbf{r})\otimes PSF(\mathbf{r})$, plus the offset (pedestal), $\langle{B_{k}}\rangle\bar{I}$, via a linear combination of speckle (noise) maps, $\langle B_{k}I_{k}(\mathbf{r})\rangle$.
@@ III Diffraction catastrophes as a random basis for ghost projection
In this paper, we propose an alternative basis for ghost projection, replacing the ensemble of speckle maps with an ensemble of focal-field diffraction catastrophes [7]. [^4] If the aberrations associated with the distorted focal field are sufficiently strong, the resulting focal-plane caustics constitute the “diffraction catastrophes” [7, 10] referred to in this sentence from the main text. If the aberrations are sufficiently weak—while also being large enough to split a focal spot into a relatively large number of local intensity maxima—the peaked focal-field caustic intensity structure will be significantly softened. In this latter case, one will instead have distorted arrays of diffraction spots such as those illustrated on p. 541 of Born and Wolf [8]. The methods of our paper may also be applied to such a case, namely where the aberrated focal-field patterns are highly structured, but the aberrations are not sufficiently strong for focal-plane caustics to be evident. A similar statement holds for a complementary limiting case, where the geometrical-optics limit implies that caustic-laden focal-field intensities may be described using ray optics. This corresponds to modifying the scenario in Fig. 2 to that shown in Fig. 1(b). The key changes are that the correlation length $\ell$ is now a function of transverse position $\mathbf{r}$ and the pedestal, $\mathcal{P}(\mathbf{r})$, is no longer spatially uniform. Equation (8) thereby generalizes to
$\left[\iint S(\mathbf{r}^{\prime})PSF(\mathbf{r},\mathbf{r}^{\prime})\,d\mathbf{r}^{\prime}\right]+\mathcal{P}(\mathbf{r})=\langle B_{k}I_{k}(\mathbf{r})\rangle,$ (9)
@@
whereby the space-invariant point-spread function $PSF(\mathbf{r})$ has been replaced with a space-variant smearing function (scaled autocovariance) $PSF(\mathbf{r},\mathbf{r}^{\prime})$ that converts the convolution into a linear integral transform, and $\{I_{k}(\mathbf{r})\}$ now denotes an ensemble of diffraction catastrophes. In practice, the position dependence of the smearing function $PSF(\mathbf{r},\mathbf{r}^{\prime})$ will often be fairly weak, if one restricts the field of view for the desired ghost projection to be in a sufficiently small vicinity of the optical axis.
@@
From this point we will utilize a vector-space description of pixellated images to lay the foundation for understanding how ghost projection functions in practice, and how it can be optimized. Notions of superposition and weighting of patterns naturally lead to treating the illuminating patterns as elements of a vector space [22], a formalism which is amenable to simulation and creating weight-generating algorithms. In ghost projection, the set of illuminating patterns can be considered as a fixed set of function-space vectors; the weight-generating aspect of ghost projection is analogous to finding a linear combination of these vectors that approximates some desired vector (image), with the constraint that the coefficients be non-negative because negative exposure times are non-physical. A consequence of this non-negativity constraint is that the image can only be synthesized up to the previously mentioned offset that we refer to as the “pedestal” [1, 2].
@@
Let the desired ghost-projection image be $S$, and the set of weights for the illuminating patterns be $\{w_{k}\}_{k\in\{1,\cdots,\mathcal{N}\}}$, where each $w_{k}\geq 0$ and each $I_{k}$ has only non-negative entries. Note that for ease of notation, we have suppressed the functional dependence on $\mathbf{r}$. The goal of the computational side of ghost projection is to find some optimum set of weights, $w_{k}$, such that in combination they provide an accurate synthesis of $S$ [1, 2, 3], i.e.,
$\sum_{k=1}^{\mathcal{N}}w_{k}I_{k}\approx S+\mathcal{P}.$ (10)
@@
    A significantly more efficient method of synthesizing the desired image, $S$, is to use a suitable iterative numerical method such as the non-negative least squares (NNLS) algorithm [2, 3]. We recall the previously mentioned fact that the weighting coefficients $w_{k}$ cannot be negative, since they are proportional to the exposure times for each element of the random basis. NNLS can be employed to determine the set $\{w_{k}\geq 0\}$ that minimizes the cost function
    $\mathcal{C}=\left\lVert\sum_{k=0}^{\mathcal{N}}w_{k}I_{k}-S\right\rVert_{2},$ (13)
@@
The optimal solution to Eq. (13) will choose a minimal set of basis patterns to avoid adding a pedestal to the result that has a square-error cost. Depending on the size of the basis set, the constructed image may not have a high quality in terms of reproducing the target image details. Should the application allow it, we can improve image fidelity by allowing a flat pedestal to be added. An effective method to achieve that is to determine the non-negative weights $w_{k}$ that minimize
$\mathcal{C}=\left\lVert\sum_{k=0}^{\mathcal{N}}w_{k}I^{\prime}_{k}-S^{\prime}\right\rVert_{2},$ (14)
where $I_{k}^{\prime}(\mathbf{r}):=I_{k}(\mathbf{r})-\overline{I_{k}}$ and $S^{\prime}(\mathbf{r}):=S(\mathbf{r})-\overline{S}$ are mean-corrected forms of $I_{k}$ and $S$. This is ideal when the ensemble-averaged pattern $\langle I(\mathbf{r})\rangle$ (henceforth abbreviated as the AIP for average illuminating pattern) is relatively flat. Further modifications of this optimization are required when the AIP is not flat (as is the case with diffraction catastrophes). This will be developed in Sec. V.3.
While approach (ii) enforces sparsity in an explicit manner, approach (iii) implicitly imposes sparsity, since NNLS typically assigns significant weights $w_{k}$ to only a small fraction of the random basis [24] (although this does depend on the target image). As a result, if we discard weights below a specified threshold, the number of patterns projected can be greatly reduced with minimal detriment to reconstruction quality. The larger the number of members in the random basis from which one can choose, the more efficient the ghost-projection process can become. [^5] We have formalized several schemes to select the sparse set of non-zero pattern weightings to produce a desired image, given an initial large set of basis patterns. These schemes will be described later in the paper.